Use properties of logarithms to expand the logarithmic expression as much as possible. Where possible, evaluate logarithmic expressions without using a calculator, ln((e^3)/5)
Answer by Alan3354(69443)

Use properties of logarithms to expand the logarithmic expression as much as possible. Where possible, evaluate logarithmic expressions without using a calculator, ln((e^3)/5)
----------------------
ln(e^3) = 3
3/5 = 3/5